Suppose a sample of 800 people is drawn. Of these people, 368 passed out. Using the data, construct the 98% confidence interval for the population proportion of people who black out at G forces greater than 6. Round your answers to three decimal places.

Lower Point: ?

Upper Point: ?
Answer #1

Solution :

Given that,

Point estimate = sample proportion = = x / n = 368 / 800 = 0.460

1 - = 1 - 0.460 = 0.54

Z/2 = 2.326

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 2.326 * (((0.460 * 0.54) / 800)

Margin of error = E = 0.041

A 98% confidence interval for population proportion p is ,

- E < p < + E

0.460 - 0.041 < p < 0.460 + 0.041

0.419 < p < 0.501

Lower Point: 0.419

Upper Point: 0.501